OpenStudy (anonymous):

The demand and supply equationsf or an electronico rganizear reg ivenb y P : 1 8 0 - 4 x p : 7 5 + 3 x Demand equation Supply equation where p is the price in dollars and x represents the number of units, in thousands. Find the equilibrium point for this market.

OpenStudy (anonymous):

hello few once again we came to answer same question what a coincidence

OpenStudy (anonymous):

How bizarre! What are the odds?!

OpenStudy (anonymous):

LADAN in short u just ask u want to find the Break Even Point

OpenStudy (anonymous):

that is not odd actually PPL are short of questions so they are posting the same questions a number of times

OpenStudy (anonymous):

leave all these come let us go to help those who really need help !

OpenStudy (anonymous):

Haha you're right.

OpenStudy (anonymous):

sorry I really need your help

OpenStudy (anonymous):

yeah i m ready to but i had answered this question dint i ??

OpenStudy (anonymous):

this question is for my major that is accounting

OpenStudy (anonymous):

I didn't see the answer.because of that I ask it again

OpenStudy (anonymous):

see here u need to calculate the Break Even Point Right. so u have to equate the demand and supply

OpenStudy (anonymous):

in terms of price the demand and supply function is given so at what point it wud be equal??

OpenStudy (anonymous):

is the answer of this question is 15??

OpenStudy (anonymous):

I think that.I just didn't know how to find a solution

OpenStudy (anonymous):

few tell me if u will equate two equation what u will do

OpenStudy (anonymous):

I can solve it

OpenStudy (anonymous):

Demand Function \[P = 1 8 0 - 4 x \] Supply Function \[P = 7 5 + 3 x \]

OpenStudy (anonymous):

how u will solve tell

OpenStudy (anonymous):

I do this. 75+3x=180-4x

OpenStudy (anonymous):

gr8 buddy u got the point

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

then 7x=105

OpenStudy (anonymous):

answer is 15

OpenStudy (anonymous):

and x=15

OpenStudy (anonymous):

I really thank you